OpenStudy (anonymous):

find the y intercept for -2x+ 2y = 4

OpenStudy (saifoo.khan):

when y intercpts, x = 0 so: -2x+ 2y = 4 -2(0)+ 2y = 4 2y = 4 y = 4/2 y = 2

OpenStudy (saifoo.khan):

(0,2)